OpenStudy (anonymous):

a^2 – a – 20 can you help me solve this idk why I keep getting messed up on this

OpenStudy (mathstudent55):

Set up two sets of parentheses: ( )( ) Place a on the left of each parentheses: (a )(a ) Your middle term is -a which means -1a, and your last term is -20 Now look for two numbers that multiply to -20 and add to -1 What are they? That's what goes inside the parenteses on the right sides.

OpenStudy (anonymous):

-5, and -4

OpenStudy (anonymous):

K so we have to factor a^2-a-20. We need to find two factors that when multiplied will equal -1 and when added equal -20. If we go through the factors of -20, we find that the two numbers that do this are -5 & 4. -5*4=-20 & -5+4=-1 so it fits what we're looking for. Now we put this in the two sets of parenthesis we have- (a )(a ) and it becomes (a+4)(a-5)

OpenStudy (anonymous):

wait y is one negative and one positive

OpenStudy (anonymous):

Because we are looking for 2 numbers that multiply to get -20. A positive*positive=positive negative*negative=positive positive*negative=negative. So that means one of the numbers will be negative & the other positive. Also these 2 numbers have to add up to -1. The only two numbers that will meet this are -5 & 4. -5*4=-20 and -5+4=-1.

OpenStudy (mathstudent55):

It's -5 and 4 because -5 x 4 = -20 and -5 + 4 = -1 So now those numbers go in the parentheses. (a - 5)(a + 4) And you're done.

OpenStudy (anonymous):

o ok thank you

OpenStudy (anonymous):

no prob. :)